A 3.00-kg model airplane has velocity components of 5.00 m/s due east and 8.00 m/s due north. What is the plane’s kinetic energy?

a) 96 J

b) 254 J

c) 38 J

d)134 J
Answer #1

Mass of plane = 3 kg

Velocity of plane = Vxi + Vyj

V = 5 i + 8 j

Magnitude of velocity = |V| = sqrt (5^2 + 8^2)

|V| = 9.43 m/sec

Kinetic energy is given by:

KE = 0.5*m*V^2

KE = 0.5*3*9.43^2 = 133.39 J

Correct option is D (134 J)